# TK-Planes: Tiered K-Planes with High Dimensional Feature Vectors for Dynamic UAV-based Scenes

### v1

@@ V Conclusion and Future Work
We have presented a novel method for synthetic data generation. Our tiered K-Planes algorithm has shown that it can improve the fidelity of dynamic objects within difficult to render scenes such as those from UAVs. In certain cases this improvement is upwards of 3 points in Dynamic-PSNR. Although there are failure cases in some instances, this can easily be explained by a lack of proper hyperparameter grid search with respect to patch size, feature vector dimension and grid resolution, amongst other parameters. Overall, the increase in fidelity of dynamic objects will help with tasks such as object detection, pose recognition and action recognition wherein the details of the subjects are paramount.
Future work with respect to TK-Planes involves a more robust grid search of hyperparameters given particular scenes and environmental conditions. Furthermore, the use of abstract feature vectors can be exploited to take advantage of redundant information within a scene and improve to a greater degree the fidelity of images and novelty of poses and time-slices of dynamic scenes. It may be useful to design some hybrid methods that can combine Extended K-Planes and TK-Planes to further improve the recognition accuracy.

### v2

@@ V Conclusion and Future Work
We have presented a novel method for synthetic data generation. Our tiered K-Planes algorithm has shown that it can improve the fidelity of dynamic objects within difficult to render scenes such as those from UAVs. In certain cases this improvement is upwards of 3.96 PSNR and 4.1 DPSNR over previous NeRF methods and 1.05 PSNR and 4.58 DPSNR over a state of the art 4D Gaussian Splatting method. Overall, the increase in fidelity of dynamic objects will help with tasks such as object detection, pose recognition and action recognition wherein the details of the subjects are paramount.
Future work with respect to TK-Planes involves a more robust grid search of hyperparameters given particular scenes and environmental conditions. Furthermore, the use of abstract feature vectors can be exploited to take advantage of redundant information within a scene and improve to a greater degree the fidelity of images and novelty of poses and time-slices of dynamic scenes.